# The tensor product of $p$-adic Hilbert spaces

### v1

@@ 1 Introduction
According to the models proposed so far, the field $\mathbb{Q}_{p}$ can be invoked at two different levels when formulating a $p$-adic quantum mechanical theory. On the one hand, one may simply resort to wave functions of the form $\psi\colon\mathbb{Q}_{p}\to\mathbb{C}$ (or, more generally, complex wave functions on a $p$-adic manifold), in such a way to incorporate the non-Archimedean behavior of the theory directly into the space-time coordinates [8, 9, 10, 11, 12, 13, 14, 15]. On the other hand, $\mathbb{Q}_{p}$ — or rather, for technical reasons related, e.g., to the role played by a suitable conjugation map in the theory, a quadratic extension of this field — may be adopted as the base field of the carrier vector space of physical states [16, 17, 18, 19, 20, 21]. This is both a more subtle and more significant modification of the standard formulation of quantum mechanics that entails, among other aspects, the development of new mathematical tools and a suitable interpretation of the physically observable quantities [22, 23, 24].
In a preceding paper [22], a new approach to quantum mechanics over the non-Archimedean field of $p$-adic numbers has been proposed that mimics, to some extent, the algebraic formulation of standard quantum mechanics [25]. The main novelty of this approach lies in a suitable definition of a $p$-adic Hilbert space and its remarkable consequences. A central aspect of our definition of a $p$-adic Hilbert space is that the existence of orthonormal bases is part of the fundamental axioms rather than an effect of these axioms. This peculiar fact is due to the very special interplay between the non-Archimedean norm and the scalar product [22], that has no parallel in the standard complex case.
As previously mentioned, in our formulation of $p$-adic quantum mechanics, physical states are defined as linear functionals on a *-algebra of observables over a quadratic extension $\mathbb{Q}_{p,\mu}$ of the field of $p$-adic numbers, where $\mu$ is a parameter labeling the various possible quadratic extensions of $\mathbb{Q}_{p}$ (for generalities regarding the quadratic extensions of $p$-adic numbers, see [6, 7]). At this stage, as pointed out in [22], an important step forward in developing a complete theory of $p$-adic quantum mechanics is to define the tensor product of two $p$-adic Hilbert spaces $\mathcal{H}$ and $\mathcal{K}$, to provide a mathematical description of a composite physical system. In our present contribution, we address this issue precisely.
@@
Finally, we study how the tensor product $\mathcal{H}\otimes\mathcal{K}$ behaves w.r.t. the subspaces of the original $p$-adic Hilbert spaces $\mathcal{H}$ and $\mathcal{K}$. In this regard, among many other technical issues, a fundamental point that one should keep in mind is the following. The notion of subspace itself is not as simple and straightforward in the $p$-adic setting as it is in the complex case [23]. Accordingly, the analogies — as well as several nontrivial differences — with respect to the standard complex setting, are analyzed throughout the article.
@@ 2.1 $p$-adic Hilbert spaces
We say that $\|\cdot\|$ is an ultrametric seminorm if it is absolute-homogeneous and satisfies (1). If $X$ is complete w.r.t. the (ultra)-metric induced by $\|\cdot\|$, the pair $(X,\|\cdot\|)$ is called a $p$-adic Banach space [7, 22, 27, 23, 29, 30, 31]. In the following, we will consider separable $p$-adic Banach spaces.
@@
holds for every $\alpha,\beta\in\mathbb{Q}_{p,\mu}$. We say that $\{x,y\}\subset X\setminus\{0\}$ is a norm-orthogonal system if $x$ and $y$ are mutually norm-orthogonal; we say that $\{x,y\}$ is a normal system if $x$ and $y$ are mutually norm-orthogonal, and, in addition, they are normalized, i.e., $\|x\|=\|y\|=1$.
@@
It is worth observing that in an inner-product $p$-adic Banach space, in general, $\|x\|\neq\sqrt{\left|\langle x,x\rangle\right|}$, i.e., the ultrametric norm does not stem directly from the non-Archimedean inner-product [22, 23]. Accordingly, in the $p$-adic setting, we have two distinct natural notions of orthogonality: the IP-orthogonality and the (previously introduced) norm-orthogonality.
A countable set of vectors $\Psi\equiv\{\psi_{i}\}_{i\in I}$ in an inner product $p$-adic Banach space $(X,\|\cdot\|)$ is said to be an orthonormal basis, if $\Psi$ is a normal basis in $X$ (see Definition 2.3) and if, moreover, $\langle\psi_{i},\psi_{j}\rangle=\delta_{ij}$, for every pair of indices $i,j\in I$. In an inner product $p$-adic Banach space, the existence of an orthonormal basis is not guaranteed. Nevertheless, when such a basis exists, several salient properties of complex Hilbert spaces have analogs in the p-adic setting. It is precisely this observation that led us to set the following [22, 23]:
@@
$c_{0}(I,\mathbb{Q}_{p,\mu})\times c_{0}(I,\mathbb{Q}_{p,\mu})\ni(x,y)\mapsto\langle x,y\rangle_{c}\coloneqq\sum_{i\in I}\overline{x}_{i}y_{i}\in\mathbb{Q}_{p,\mu}.$ (13)
By construction, we have: $\langle e_{i},e_{j}\rangle_{c}=\delta_{ij}$, $\forall i,j\in I$, i.e., $\bm{e}\equiv\{e_{i}\}_{i\in I}$ is an orthonormal basis. Therefore, $(c_{0}(I,\mathbb{Q}_{p,\mu}),\|\cdot\|_{\infty},\langle\hskip 0.9105pt\cdot\hskip 1.70717pt,\cdot\hskip 0.9105pt\rangle_{c},\bm{e})$ is a $p$-adic Hilbert space; in the literature [22, 23, 34, 35] it is sometimes referred to as the coordinate $p$-adic Hilbert space, and denoted by $\mathbb{H}(I)$.
@@
$\ell^{\infty}(I,\mathbb{Q}_{p,\mu})\coloneqq\big\{\eta\equiv\{\eta_{i}\}_{i\in I}\mid\eta_{i}\in\mathbb{Q}_{p,\mu},\,\eta\,\mbox{bounded sequence in $\mathbb{Q}_{p,\mu}$}\big\}.$ (15)
@@
$\mathcal{L}_{\Psi}\colon\ell^{\infty}(I,\mathbb{Q}_{p,\mu})\ni\xi=\{\xi_{i}\}_{i\in I}\mapsto\sum_{i\in I}\xi_{i}\langle\psi_{i},\cdot\hskip 1.13809pt\rangle=\sum_{i\in I}\xi_{i}\mathcal{J}_{\mbox{\tiny$\mathcal{H}$}}\psi_{i}\in\mathcal{H}^{\prime},$ (17)
@@ 2.2.1 Bounded and adjointable operators
As in the complex setting, $B\colon\mathcal{H}\dashrightarrow\mathcal{K}$ is bounded iff it is continuous [22, 23]; moreover, if $B$ is bounded, it can be extended by continuity to the norm-closure of its domain $\mathrm{dom}(B)$, which is a $p$-adic Banach space, but, in general, not a $p$-adic Hilbert space. Therefore, in the $p$-adic setting (bounded) operators whose domain coincides with the full initial space deserve special attention, and we adopt a ad hoc terminology and notation.
@@
$\mathcal{B}(\mathcal{H},\mathcal{K})=\big\{\mathrm{op}_{\Phi,\Psi}(B_{mn})\colon\mathcal{H}\dashrightarrow\mathcal{K}\mid\mbox{$(B_{mn})\in\mathsf{M}_{\infty}(\mathbb{Q}_{p,\mu})$, $\mathrm{dom}(\mathrm{op}_{\Phi,\Psi}(B_{mn}))=\mathcal{H}$}\big\}=(\mathcal{H},\mathcal{K})_{\Phi,\Psi}.$ (25)
@@
$\mathrm{dom}\big(B^{\dagger}\big)\coloneqq\{\psi\in\mathcal{K}\mid\mbox{there exists $\zeta\in\mathcal{H}$ s.t.\ $\langle\psi,B\phi\rangle=\langle\zeta,\phi\rangle$, $\forall\phi\in\mathcal{H}=\mathrm{dom}(B)$}\}.$ (28)
@@
In what follows, we will denote by $\mathcal{B}_{\mathrm{ad}}(\mathcal{H},\mathcal{K})$ the set of all adjointable bounded operators from $\mathcal{H}$ to $\mathcal{K}$. We now aim to provide a suitable matrix characterization of an adjointable operator. In particular, we prove the following:
@@ 2.2.2 The $p$-adic trace/Hilbert-Schmidt class
Moreover, for every pair of orthonormal bases $\Phi\equiv\{\phi_{m}\}_{m\in\mathbb{N}}$ in $\mathcal{H}$ and $\Psi\equiv\{\psi_{n}\}_{n\in\mathbb{N}}$ in $\mathcal{K}$, $\mathcal{T}(\mathcal{H},\mathcal{K})$ coincides with the uniform closure of the set $\mathscr{B}_{\Phi,\Psi}$ of block-finite operators w.r.r.t the orthonormal bases $\Phi$, $\Psi$; i.e.,
@@
The first claim is proved likewise Theorem $6.35$ in [22], whereas the second is a straightforward adaptation of Theorem $6.34$ in [22] to our specific case. ∎
@@
$\hskip-1.99168pt\sideset{{}^{jk}\hskip-1.42262pt}{{}^{\Phi,\Psi}}{\operatorname{\mathnormal{E}}}\coloneqq\mathrm{op}_{\Phi,\Psi}\big(\hskip-1.99168pt\sideset{{}^{jk}\hskip-1.42262pt}{{}_{mn}}{\operatorname{\mathnormal{E}}}\big),\quad\mbox{where $\hskip-1.99168pt\sideset{{}^{jk}\hskip-1.42262pt}{{}_{mn}}{\operatorname{\mathnormal{E}}}=\delta_{jm}\delta_{kn}$};$ (43)
i.e., in Dirac’s notation, we have: $\hskip-1.99168pt\sideset{{}^{jk}\hskip-1.42262pt}{{}^{\Phi,\Psi}}{\operatorname{\mathnormal{E}}}=|\psi_{j}\rangle\langle\phi_{k}|\equiv\langle\phi_{k},\cdot\hskip 1.13809pt\rangle\psi_{j}$. Note that, for every $j,k\in\mathbb{N}$, $\hskip-1.99168pt\sideset{{}^{jk}\hskip-1.42262pt}{{}^{\Phi,\Psi}}{\operatorname{\mathnormal{E}}}$, is a rank-one operator and, therefore, it is in the trace class $\mathcal{T}(\mathcal{H},\mathcal{K})$.
@@ 3 Construction of the tensor product
for all vectors $x,x_{1},x_{2}\in X$, $y,y_{1},y_{2}\in Y$, and all scalars $\alpha_{i},\beta_{i}\in\mathbb{Q}_{p,\mu}$, $i=1,2$. We will denote by $\mathsf{Bil}\hskip 0.85358pt(X\times Y,Z)$ the vector space of all bilinear maps from $X\times Y$ to $Z.$ In the special case where $Z=\mathbb{Q}_{p,\mu}$ (the one-dimensional $\mathbb{Q}_{p,\mu}$-linear space), we put $\mathsf{Bil}\hskip 0.85358pt(X\times Y)\equiv\mathsf{Bil}\hskip 0.85358pt(X\times Y,\mathbb{Q}_{p,\mu})$ (the space of bilinear forms on $X\times Y$).
@@
For every bilinear map $\hskip 0.56905pt\mathfrak{B}\hskip 0.56905pt:X\times Y\to Z$, there exists a unique linear map $B\colon X\otimes_{\alpha}Y\to Z$ such that $\hskip 0.56905pt\mathfrak{B}\hskip 0.56905pt(x,y)=B(x\otimes y)$, for all $x\in X$ and $y\in Y$.
@@
Let $\hskip 0.56905pt\mathfrak{B}\hskip 0.56905pt:X\times Y\to Z$ be a bilinear map and define $B:X\otimes_{\alpha}Y\to Z$ by
@@
Suppose now that $\sum_{i=1}^{n}x_{i}\otimes y_{i}=0.$ Then, for each linear functional $\varphi$ on $Z,$ the composition $\varphi\circ B$ is a bilinear functional on $X\times Y$ and
@@
Therefore, $B$ is well defined and the bilinear map $\hskip 0.56905pt\mathfrak{B}\hskip 0.56905pt$ associated with the linear map $B$ yields an identification of $\mathsf{Bil}\hskip 0.85358pt(X\times Y,Z)$ and linear maps from $X\otimes_{\alpha}Y$ to Z. ∎
@@
    if $x_{1},\dots,x_{n}\in X$ and $y_{1},\dots,y_{n}\in Y$ — with $y_{1},\dots,y_{n}$ linearly independent and $x_{i}\neq 0$, for all $i\in\{1,\ldots,n\}$ — then $\{x_{1}\otimes y_{1},\dots,x_{n}\otimes y_{n}\}$ is linearly independent set.
@@
In addition to the properties (i)(vi) listed in Proposition 3.4, we have the following further useful fact, namely, a characterization of the zero vector in $X\otimes_{\alpha}Y$.
@@
Let $X,Y$ be vector spaces over $\mathbb{Q}_{p,\mu}$, and let $X^{\vee},Y^{\vee}$ the associated algebraic dual spaces. Then, the following statements are equivalent for $u=\sum_{i=1}^{n}x_{i}\otimes y_{i}\in X\otimes_{\alpha}Y$:
@@
Therefore, the next step of our construction is to endow the algebraic tensor product $\mathcal{H}\otimes_{\alpha}\mathcal{K}$ with a suitable ultrametric norm, in such a way as to obtain, upon completion, a suitable $p$-adic Banach space.
Recall [40] that in the standard setting where the tensor product $\mathscr{H}\otimes\mathscr{K}$ of two complex Hilbert spaces $\mathscr{H}$, $\mathscr{K}$ is considered, the norm of the Hilbert space $\mathscr{H}\otimes\mathscr{K}$ is precisely the one canonically associated with the scalar product of this space. Otherwise stated, a suitable scalar product is defined on the algebraic tensor product $\mathscr{H}\otimes_{\alpha}\mathscr{K}$, so obtaining a pre-Hilbert space, and the (Hilbert) tensor product $\mathscr{H}\otimes\mathscr{K}$ is then achieved as the Hilbert space completion of this space. This approach cannot be pursued in the $p$-adic setting, because, as noted previously, the norm and the scalar product are not related in the usual fashion — but only by the Cauchy-Schwarz inequality — in this case. It is also worth noting that the canonical (Hilbert space) norm on $\mathscr{H}\otimes\mathscr{K}$ does not coincide with the projective norm [26] defined on the algebraic tensor product $\mathscr{H}\otimes_{\alpha}\mathscr{K}$ (the linear spaces $\mathscr{H}$, $\mathscr{K}$ being endowed with their respective Hilbert space norms); namely, the (Hilbert space) norm on $\mathscr{H}\otimes\mathscr{K}$ does not coincide with the following:
@@
Here, the infimum is taken over all possible representations of $u$.
@@
Here, the infimum is taken, once again, over all possible representations of $u\in\mathcal{H}\otimes_{\alpha}\mathcal{K}$.
We now demonstrate that it indeed defines a non-Archimedean norm. To this end, we first prove that the mapping $\mathcal{H}\otimes_{\alpha}\mathcal{K}\ni u\mapsto\|u\|_{\pi}\in\mathbb{R^{+}}$ in (54) is an ultrametric seminorm on the algebraic tensor product $\mathcal{H}\otimes_{\alpha}\mathcal{K}.$
@@
Starting with linearity, we have to prove that $\|\lambda u\|_{\pi}=\left|\lambda\right|\|u\|_{\pi}$. For $\lambda=0$, this is trivial. If $\lambda\neq 0,$ then $\lambda u=\sum_{i=1}^{n}(\lambda x_{i})\otimes y_{i}$ is a possible representation of $\lambda u$ and, hence, we have:
@@
On the other hand we can write $\|u\|_{\pi}=\|\lambda^{-1}\lambda u\|_{\pi}$ which entails
@@
Thus from (55) and (56) we have $\|\lambda u\|_{\pi}=\left|\lambda\right|\|u\|_{\pi}.$ It remains to prove the $\|\cdot\|_{\pi}$ strong triangular inequality. For $\epsilon>0$ and let $u=\sum_{i=1}^{n}x_{i}\otimes y_{i}$ be a possible representation of $u$ and $v=\sum_{i=n+1}^{n+m}x_{i}\otimes y_{i}$ be a possible representation of $v$ such that

### v2

@@ 1 Introduction
According to the models proposed so far, the field $\mathbb{Q}_{p}$ can be invoked at two different levels when formulating a $p$-adic quantum mechanical theory. On the one hand, one may simply resort to wave functions of the form $\psi\colon\mathbb{Q}_{p}\to\mathbb{C}$ (or, more generally, complex wave functions on a $p$-adic manifold), in such a way as to incorporate the non-Archimedean behavior of the theory directly into the space-time coordinates [8, 9, 10, 11, 12, 13, 14, 15]. On the other hand, $\mathbb{Q}_{p}$ — or rather, for technical reasons related, e.g., to the role played by a suitable conjugation map in the theory, a quadratic extension of this field — may be adopted as the base field of the carrier vector space of physical states [16, 17, 18, 19, 20, 21]. This is both a more subtle and more significant modification of the standard formulation of quantum mechanics that involves, among other aspects, the development of new mathematical tools and a suitable interpretation of the physically observable quantities [22, 23, 24].
In a preceding paper [22], a new approach to quantum mechanics over the non-Archimedean field of $p$-adic numbers has been proposed that mimics, to some extent, the algebraic formulation of standard quantum mechanics [25]. The main novelty of this approach lies in a suitable definition of a $p$-adic Hilbert space and its remarkable consequences. A central aspect of our definition of a $p$-adic Hilbert space is that the existence of orthonormal bases is part of the fundamental axioms rather than an effect of these axioms. This peculiar fact is due to the very special interplay between the non-Archimedean norm and the scalar product [22], which has no parallel in the standard complex case.
As mentioned previously, in our formulation of $p$-adic quantum mechanics, physical states are defined as linear functionals on a *-algebra of observables over a quadratic extension $\mathbb{Q}_{p,\mu}$ of the field of $p$-adic numbers, where $\mu$ is a parameter labeling the various possible quadratic extensions of $\mathbb{Q}_{p}$ (for generalities regarding the quadratic extensions of $p$-adic numbers, see [6, 7]). At this stage, as pointed out in [22], an important step forward in developing a complete theory of $p$-adic quantum mechanics is to define the tensor product of two $p$-adic Hilbert spaces $\mathcal{H}$ and $\mathcal{K}$, to provide a mathematical description of a composite physical system. In our present contribution, we address this issue precisely.
@@
Finally, we study how the tensor product $\mathcal{H}\otimes\mathcal{K}$ behaves w.r.t. the subspaces of the original $p$-adic Hilbert spaces $\mathcal{H}$ and $\mathcal{K}$. In this regard, among many other technical issues, a fundamental point that should be considered is the following. The notion of subspace itself is not as simple and straightforward in the $p$-adic setting as it is in the complex case [23]. Accordingly, the analogies — as well as several nontrivial differences — with respect to the standard complex setting, are analyzed throughout the article.
@@ 2.1 $p$-adic Hilbert spaces
(More generally, we say that $\|\cdot\|\colon X\rightarrow\mathbb{R}^{+}$ is an ultrametric seminorm if it is absolute-homogeneous and satisfies relation (1).) If $X$ is complete w.r.t. the (ultra)-metric induced by the norm $\|\cdot\|$, the pair $(X,\|\cdot\|)$ is called a $p$-adic Banach space [7, 22, 27, 23, 29, 30, 31]. In the following, we will consider separable $p$-adic Banach spaces only.
@@
holds for every $\alpha,\beta\in\mathbb{Q}_{p,\mu}$. We say that $\{x,y\}\subset X\setminus\{0\}$ is a norm-orthogonal system if $x$ and $y$ are mutually norm-orthogonal; in particular, we say that $\{x,y\}$ is a normal system if $x$ and $y$ are mutually norm-orthogonal, and, in addition, they are normalized, i.e., $\|x\|=\|y\|=1$.
@@
If $\mathfrak{B}$ is a norm-orthogonal basis in a $p$-adic Banach space $(X,\|\cdot\|)$, then, for every vector $x\in X$, we have [22]:
$\|x\|=\sup_{b\in\mathfrak{B}}|\alpha_{x}(b)|\hskip 1.70717pt\|b\|=\max_{b\in\mathfrak{B}}|\alpha_{x}(b)|\hskip 1.70717pt\|b\|.$ (4)
@@
It is worth observing that in an inner-product $p$-adic Banach space, in general, $\|x\|\neq\sqrt{\left|\langle x,x\rangle\right|}$, i.e., the ultrametric norm does not stem directly from the non-Archimedean inner product [22, 23]. Accordingly, in the $p$-adic setting, we have two distinct natural notions of orthogonality: the IP-orthogonality and the (previously introduced) norm-orthogonality.
A countable set of vectors $\Psi\equiv\{\psi_{i}\}_{i\in I}$ in an inner-product $p$-adic Banach space $(X,\|\cdot\|)$ is said to be an orthonormal basis, if $\Psi$ is a normal basis in $X$ (see Definition 2.3) and if, moreover, $\langle\psi_{i},\psi_{j}\rangle=\delta_{ij}$, for every pair of indices $i,j\in I$. In an inner product $p$-adic Banach space, the existence of an orthonormal basis is not guaranteed. Nevertheless, when such a basis exists, several salient properties of complex Hilbert spaces have analogs in the p-adic setting. It is precisely this observation that led us to set the following [22, 23]:
@@
$c_{0}(I,\mathbb{Q}_{p,\mu})\times c_{0}(I,\mathbb{Q}_{p,\mu})\ni(x,y)\mapsto\langle x,y\rangle_{c}\coloneqq\sum_{i\in I}\overline{x}_{i}y_{i}\in\mathbb{Q}_{p,\mu},$ (14)
where the overline denotes the conjugate in $\mathbb{Q}_{p,\mu}$. By construction, we have: $\langle e_{i},e_{j}\rangle_{c}=\delta_{ij}$, $\forall i,j\in I$, i.e., $\bm{e}\equiv\{e_{i}\}_{i\in I}$ is an orthonormal basis. Therefore, $(c_{0}(I,\mathbb{Q}_{p,\mu}),\|\cdot\|_{\infty},\langle\hskip 0.9105pt\cdot\hskip 1.70717pt,\cdot\hskip 0.9105pt\rangle_{c},\bm{e})$ is a $p$-adic Hilbert space; in the literature [22, 23, 34, 35] it is sometimes referred to as the coordinate $p$-adic Hilbert space, and denoted by $\mathbb{H}(I)$.
@@
$\ell^{\infty}(I,\mathbb{Q}_{p,\mu})\coloneqq\big\{\eta\equiv\{\eta_{i}\}_{i\in I}\mid\eta_{i}\in\mathbb{Q}_{p,\mu},\,|\eta_{i}|<\infty\big\}.$ (16)
@@
$\mathcal{E}_{\Psi}\colon\ell^{\infty}(I,\mathbb{Q}_{p,\mu})\ni\xi=\{\xi_{i}\}_{i\in I}\mapsto\sum_{i\in I}\xi_{i}\langle\psi_{i},\cdot\hskip 1.13809pt\rangle=\sum_{i\in I}\xi_{i}\mathcal{J}_{\mbox{\tiny$\mathcal{H}$}}\psi_{i}\in\mathcal{H}^{\prime},$ (18)
@@ 2.2.1 Bounded and adjointable operators
As in the complex setting, $B\colon\mathcal{H}\dashrightarrow\mathcal{K}$ is bounded iff it is continuous [22, 23]; moreover, if $B$ is bounded, it can be extended by continuity to the norm-closure of its domain $\mathrm{dom}(B)$, which is a $p$-adic Banach space, but, in general, not a $p$-adic Hilbert space. Therefore, in the $p$-adic setting (bounded) operators whose domain coincides with the full initial space deserve special attention, and we adopt an ad hoc terminology and notation.
@@
$\mathcal{B}(\mathcal{H},\mathcal{K})$ $=(\mathcal{H},\mathcal{K})_{\Phi,\Psi}$
$=\big\{\mathrm{op}_{\Phi,\Psi}(B_{mn})\colon\mathcal{H}\dashrightarrow\mathcal{K}\mid\mbox{\rm$\lim_{m}B_{mn}=0$, $n\in\mathbb{N}$, and $\sup_{m,n}|B_{mn}|<\infty$}\big\}.$ (26)
@@
$\mathrm{dom}\big(B^{\dagger}\big)\coloneqq\{\psi\in\mathcal{K}\mid\mbox{ $\exists\,\zeta\in\mathcal{H}$ s.t.\ $\langle\psi,B\phi\rangle=\langle\zeta,\phi\rangle$, $\forall\phi\in\mathcal{H}=\mathrm{dom}(B)$}\}.$ (29)
@@
In what follows, we will denote by $\mathcal{B}_{\mathrm{ad}}(\mathcal{H},\mathcal{K})$ the set of all adjointable bounded operators from $\mathcal{H}$ to $\mathcal{K}$. We now aim to provide a suitable matrix characterization of an adjointable operator. In particular, we state the following:
@@ 2.2.2 The $p$-adic trace/Hilbert-Schmidt class
Moreover, for every pair of orthonormal bases $\Phi\equiv\{\phi_{m}\}_{m\in\mathbb{N}}$ in $\mathcal{H}$ and $\Psi\equiv\{\psi_{n}\}_{n\in\mathbb{N}}$ in $\mathcal{K}$, $\mathcal{T}(\mathcal{H},\mathcal{K})$ coincides with the uniform closure of the set $\mathscr{B}_{\Phi,\Psi}$ of block-finite operators w.r.t. the orthonormal bases $\Phi$, $\Psi$; i.e.,
@@
The first claim is proved likewise Theorem $6.35$ in [22], whereas the second is a straightforward adaptation of Theorem $6.34$ in [22] to this specific case. ∎
@@
$\hskip-1.99168pt\sideset{{}^{jk}\hskip-1.42262pt}{{}^{\Phi,\Psi}}{\operatorname{\mathnormal{E}}}\coloneqq\mathrm{op}_{\Phi,\Psi}\big(\hskip-1.99168pt\sideset{{}^{jk}\hskip-1.42262pt}{{}_{mn}}{\operatorname{\mathnormal{E}}}\big),\quad\mbox{where $\hskip-1.99168pt\sideset{{}^{jk}\hskip-1.42262pt}{{}_{mn}}{\operatorname{\mathnormal{E}}}=\delta_{jm}\delta_{kn}$},$ (44)
i.e., $\hskip-1.99168pt\sideset{{}^{jk}\hskip-1.42262pt}{{}^{\Phi,\Psi}}{\operatorname{\mathnormal{E}}}=\langle\phi_{k},\cdot\hskip 1.13809pt\rangle\psi_{j}$, or equivalently $\hskip-1.99168pt\sideset{{}^{jk}\hskip-1.42262pt}{{}^{\Phi,\Psi}}{\operatorname{\mathnormal{E}}}=|\psi_{j}\rangle\langle\phi_{k}|$ in Dirac’s notation. Note that, for every $j,k\in\mathbb{N}$, $\hskip-1.99168pt\sideset{{}^{jk}\hskip-1.42262pt}{{}^{\Phi,\Psi}}{\operatorname{\mathnormal{E}}}$, is a rank-one operator and, therefore, it is in the trace class $\mathcal{T}(\mathcal{H},\mathcal{K})$.
@@ 3 Construction of the tensor product
for all vectors $x,x_{1},x_{2}\in X$, $y,y_{1},y_{2}\in Y$, and for all scalars $\alpha_{i},\beta_{i}\in\mathbb{Q}_{p,\mu}$, $i=1,2$. We will denote by $\mathsf{Bil}\hskip 0.85358pt(X\times Y,Z)$ the vector space of all bilinear maps from $X\times Y$ to $Z.$ In the special case where $Z=\mathbb{Q}_{p,\mu}$ (the one-dimensional $\mathbb{Q}_{p,\mu}$-linear space), we put $\mathsf{Bil}\hskip 0.85358pt(X\times Y)\equiv\mathsf{Bil}\hskip 0.85358pt(X\times Y,\mathbb{Q}_{p,\mu})$ (the space of bilinear forms on $X\times Y$).
@@
For every bilinear map $\hskip 0.56905pt\mathfrak{B}\hskip 0.56905pt:X\times Y\to\mathbb{Q}_{p,\mu}$, there exists a unique linear map $B\colon X\otimes_{\alpha}Y\to\mathbb{Q}_{p,\mu}$ such that $\hskip 0.56905pt\mathfrak{B}\hskip 0.56905pt(x,y)=B(x\otimes y)$, for all $x\in X$ and $y\in Y$.
@@
Let $\hskip 0.56905pt\mathfrak{B}\hskip 0.56905pt:X\times Y\to\mathbb{Q}_{p,\mu}$ be a bilinear map and define $B:X\otimes_{\alpha}Y\to\mathbb{Q}_{p,\mu}$ by
@@
Suppose now that $\sum_{i=1}^{n}x_{i}\otimes y_{i}=0.$ Then, for each linear functional $\varphi$ on $\mathbb{Q}_{p,\mu}$, the composition $\varphi\circ B$ is a bilinear functional on $X\times Y$ and
@@
Therefore, $B$ is well defined and the bilinear map $\hskip 0.56905pt\mathfrak{B}\hskip 0.56905pt$ associated with the linear map $B$ yields an identification of $\mathsf{Bil}\hskip 0.85358pt(X\times Y)$ and linear maps from $X\otimes_{\alpha}Y$ to $\mathbb{Q}_{p,\mu}$. ∎
@@
    if $x_{1},\dots,x_{n}\in X$ and $y_{1},\dots,y_{n}\in Y$ — with $y_{1},\dots,y_{n}$ linearly independent and $x_{i}\neq 0$, for all $i\in\{1,\ldots,n\}$ — then $\{x_{1}\otimes y_{1},\dots,x_{n}\otimes y_{n}\}$ is a linearly independent set.
@@
In addition to the properties (i)(vi) listed in Proposition 3.4, we have the following further useful fact, namely, a characterization of the null vector in $X\otimes_{\alpha}Y$.
@@
Let $X,Y$ be vector spaces over $\mathbb{Q}_{p,\mu}$, and let $X^{\vee},Y^{\vee}$ be the associated algebraic dual spaces. Then, the following statements are equivalent for $u=\sum_{i=1}^{n}x_{i}\otimes y_{i}\in X\otimes_{\alpha}Y$:
@@
Therefore, the next step of our construction is to endow the algebraic tensor product $\mathcal{H}\otimes_{\alpha}\mathcal{K}$ with a suitable ultrametric norm, thereby obtaining, upon completion, a suitable $p$-adic Banach space.
Recall that in the standard setting where the tensor product $\mathscr{H}\otimes\mathscr{K}$ of two complex Hilbert spaces $\mathscr{H}$, $\mathscr{K}$ is considered, the norm of the Hilbert space $\mathscr{H}\otimes\mathscr{K}$ is precisely the one canonically associated with the scalar product of this space [41].
In other words, a suitable scalar product is defined on the algebraic tensor product $\mathscr{H}\otimes_{\alpha}\mathscr{K}$, yielding a pre-Hilbert space; the Hilbert tensor product $\mathscr{H}\otimes\mathscr{K}$ is then obtained by completing this space. Such an approach cannot be pursued in the $p$-adic setting, because, as noted previously, the norm and the scalar product are not related in the usual fashion — but only by the Cauchy-Schwarz inequality. It is also worth noting that the canonical (Hilbert space) norm on $\mathscr{H}\otimes\mathscr{K}$ does not coincide with the projective norm [26] on the algebraic tensor product $\mathscr{H}\otimes_{\alpha}\mathscr{K}$ (the linear spaces $\mathscr{H}$, $\mathscr{K}$ being endowed with their respective Hilbert space norms); namely, the (Hilbert space) norm of $\mathscr{H}\otimes\mathscr{K}$, when restricted to $\mathscr{H}\otimes_{\alpha}\mathscr{K}$, does not coincide with the following:
@@
Here, the infimum is taken over all possible representations of $u\in\mathscr{H}\otimes_{\alpha}\mathscr{K}$ (i.e., all possible decompositions of $u$ as a sum of elementary tensors).
@@
Here, the infimum is taken over all possible representations of $u\in\mathcal{H}\otimes_{\alpha}\mathcal{K}$.
We now demonstrate that (55) indeed defines a non-Archimedean norm. To this end, we first prove that the mapping $\mathcal{H}\otimes_{\alpha}\mathcal{K}\ni u\mapsto\|u\|_{\pi}\in\mathbb{R^{+}}$ in (55) is an ultrametric seminorm on the algebraic tensor product $\mathcal{H}\otimes_{\alpha}\mathcal{K}.$
@@
Let us start with absolute homogeneity. We need to prove that, for every scalar $\lambda$ and every vector $u=\sum_{i=1}^{n}x_{i}\otimes y_{i}\in\mathcal{H}\otimes_{\alpha}\mathcal{K}$, $\|\lambda u\|_{\pi}=\left|\lambda\right|\|u\|_{\pi}$. For $\lambda=0$, this is trivial. If $\lambda\neq 0,$ then $\lambda u=\sum_{i=1}^{n}(\lambda x_{i})\otimes y_{i}$ is a representation of $\lambda u$; therefore, we have the following:
@@
On the other hand, we can write $\|u\|_{\pi}=\|\lambda^{-1}\lambda u\|_{\pi}$, so that
@@
Thus, by (56) and (57) we actually have that $\|\lambda u\|_{\pi}=\left|\lambda\right|\|u\|_{\pi}$.
It remains to prove that $\|\cdot\|_{\pi}$ satisfies the strong triangle inequality. In fact, for $\epsilon>0$, let $u=\sum_{i=1}^{n}x_{i}\otimes y_{i}$ be a representation of $u$ and $v=\sum_{i=n+1}^{n+m}x_{i}\otimes y_{i}$ be representation of $v$ such that